OpenStudy (anonymous):

a metallurgist has an alloy with 15% titanium and an alloy with 20% titanium. He needs 100 grams of an alloy with 17% titanium. How much of each alloy should be mixed to attain the 100 grams of alloy with 71% titanium?

OpenStudy (anonymous):

this problem is weird. do you have the answer? \ i came up with an answer but it assumed that there is the same amount of both alloys

OpenStudy (anonymous):

yeah the answer came out to be 60 grams of the alloy with 10% of titanium needed and 40 grams of the alloy with 25% of titanium are needed. I don't know how that was the answer...

OpenStudy (anonymous):

sorry for being so late and for calling it weird. its actually fine. here is how to solve it: let x = grams of alloy A and y = grams of alloy B. now i just noticed you gave 2 sets of A and B. in the initial problem its 15% and 20% and just above its 10% and 25% for the initial assumption we have: x + y = 100 which is mass of A + mass of B = 100 and 0.15x + 0.20y = 0.17*100 which is fractionA*massA + fractionB*massB = fractionTotal*massTotal the first equation can be rearranged to x = 100 - y and we can plug that into the second equation to reduce it to a 1 variable equation: 0.15(100 - y) + 0.20y = 17 15 - 0.15y + 0.20y = 17 0.05y = 2 y = 40 grams x = 100 - y = 60 grams